# A multiplicity formula for the Milnor number of smoothable curves

### v2

@@ 5. Examples
Let $f_{0}=2y^{2}+xz+3yz+z^{2}-4x^{3}-7x^{2}y-2xy^{2}-7x^{2}z-2xyz-3x^{4}$ be a generic combination of the $2\times 2$ minors of the Jacobian matrix. Then (via a Singular computation [DGPS]) $e(\mathrm{Jac}(X_{0}))=\dim_{\mathbb{C}}\mathcal{O}_{X_{0},0}/(f_{0})=8.$
@@
But $W_{0}^{\prime}=(y,z)$. So $I_{0}(X_{0},W_{0}^{\prime})=3$. A smoothing of $X_{0}$ is given by the $2\times 2$ minors of $\begin{pmatrix}x&y&z\\ y&z&x^{2}+t\end{pmatrix}.$ Then
$X\times_{\mathbb{C}^{4}}W^{\prime}=\mathbb{V}(x,y,z)\cup\mathbb{V}(y,z,x^{2}+t)$
and $X_{t}$ and $W^{\prime}_{t}$ meet transversally at $3$ points for $t\neq 0$. As predicted by Gaffney’s Multiplicity-Polar Theorem (see Theorem 6.1)
@@
Finally, observe that for the special choice $Z^{\prime\prime}_{0}=\mathbb{V}(f_{1},f_{2})$, in the induced deformation of $Z^{\prime\prime}$ by the smoothing of $X_{0}$ above, we have $W^{\prime\prime}=\mathbb{V}(z,x^{2}-t)$ and $X\times_{\mathbb{C}^{4}}W^{\prime\prime}=\mathbb{V}(z,y^{2},x^{2}-t)$. For $t\neq 0$, one finds that $W_{t}^{\prime\prime}$ consists of two lines each tangent to $X_{t}$, i.e. $X_{t}$ and $W^{\prime\prime}_{t}$ do not intersect transversally.

### v3

@@ 5. Examples
Let $f_{0}=2y^{2}+xz+3yz+z^{2}-4x^{3}-7x^{2}y-2xy^{2}-7x^{2}z-2xyz-3x^{4}$. It is a generic combination of the $2\times 2$ minors of the Jacobian matrix. Then (via a Singular computation [DGPS]) $e(\mathrm{Jac}(X_{0}))=\dim_{\mathbb{C}}\mathcal{O}_{X_{0},0}/(f_{0})=8.$
@@
But $W_{0}^{\prime}=(y,z)$. So $I_{0}(X_{0},W_{0}^{\prime})=3$. A smoothing of $X_{0}$ is given by the $2\times 2$ minors of $\begin{pmatrix}x&y&z\\ y&z&x^{2}-t\end{pmatrix}.$ Then
$X\times_{\mathbb{C}^{4}}W^{\prime}=\mathbb{V}(x,y,z)\cup\mathbb{V}(y,z,x^{2}-t)$
and $X_{t}$ and $W^{\prime}_{t}$ meet locally transversally at $3$ points for $t\neq 0$. As predicted by Gaffney’s Multiplicity-Polar Theorem (see Theorem 6.1)
@@
Finally, observe that for the special choice $Z^{\prime\prime}_{0}=\mathbb{V}(f_{1},f_{2})$, in the induced deformation of $Z^{\prime\prime}$ by the smoothing of $X_{0}$ above, we have $W^{\prime\prime}=\mathbb{V}(z,x^{2}-t)$ and $X\times_{\mathbb{C}^{4}}W^{\prime\prime}=\mathbb{V}(z,y^{2},x^{2}-t)$. For $t\neq 0$, one finds that $W_{t}^{\prime\prime}$ consists of two lines each tangent to $X_{t}$, i.e. $X_{t}$ and $W^{\prime\prime}_{t}$ do not intersect locally transversally. As $t\rightarrow 0$, these two lines degenerate to the double line $W^{\prime\prime}_{0}=\mathbb{V}(z,x^{2})$.